Monday, 24 August

Monday, 24 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

Karl Sigmund, University of Vienna, Vienna, Austria The Population Dynamics of Conflict and Cooperation (I: p.21)

Chair: Frank Hoppensteadt, Arizona State University, Tempe, USA

10:15 – 11:15  Auditorium Maximum H 105

Michel Talagrand, CNRS, Université Paris 6, Paris, France Huge Random Structures and Mean Field Models for Spin Glasses (I: p.23)

Chair: Lior Tzafriri, The Hebrew University, Jerusalem, Israel

11:45 – 12:45  Auditorium Maximum H 105

Cumrun Vafa, Harvard University, Cambridge, USA and Tehran, Iran Geometric Physics (I: p.25)

Chair: Maxim Kontsevich, IHES, Bures-sur-Yvette, France

Invited Lectures

\underline{14:00 – 14:45}

02 Sergei Ivanov, University of Illinois, Urbana, USA On the Burnside problem for groups of even exponent (I: p.38) Chair: Otto H. Kegel

MA 005

05 Ulrich Pinkall, TU Berlin, Berlin, Germany Quaternionic line bundles and the differential geometry of surfaces (I: p.75) Chair: Dirk Ferus

MA 001

10 Dmitri Yafaev, Université Rennes-1, Rennes, France Scattering theory: some old and new problems (I: p.127) Chair: Volker Enß

H 1012

12 Friedrich Götze, Universität Bielefeld, Bielefeld, Germany Lattice Point Problems and the Central Limit Theorem in Euclidean Spaces (I: p.143) Chair: Richard M. Dudley

H 1028

13 András Frank, Eötvös Lorand University, Budapest, Hungary Applications of Relaxed Submodularity (I: p.155) Chair: Alexander Schrijver

H 2032

19 Jeremy John Gray, The Open University, Milton Keynes, England The Riemann-Roch Theorem and Geometry, 1854-1914 (I: p.208) Chair: Umberto Bottazzini

MA 004

\underline{15:00 – 15:45}

02 Idun Reiten, Norwegian University of Science and Technology, Trondheim, Norway Tilting Theory and Quasitilted algebras (I: p.43) Chair: Claus Michael Ringel

MA 005

05 Tobias Holck Colding, Courant Institute, New York University, New York, USA Manifolds and spaces with Ricci curvature bounds (I: p.68) Chair: Dirk Ferus

MA 001

10 Hans Lindblad, University of California at San Diego, La Jolla, USA Minimal regularity solutions of nonlinear wave equations (I: p.122) Chair: Konrad Gröger H 1012 12 & 16 Marco Avellaneda, Courant Institute, New York University, New York, USA Inverse Problems in Quantitative Financial Modeling (I: p.140) Chair: George C. Papanicolaou

H 1028

13 Jiří Matoušek, Charles University, Praha, Czech Republic Mathematical snapshots from the computational geometry landscape (I: p.158) Chair: Jaroslav Nešetril

H 2032

19 Karine Chemla, CNRS and Université Paris 7, Paris, France History of Mathematics in China: A factor in World History and a Source for New Questions (I: p.206) Chair: Sergei S. Demidov

MA 004

\underline{16:15 – 17:00}

03 François Gramain, Université Jean Monnet Saint-Etienne, Saint-Etienne, France Some results of algebraic independence (I: p.51) Chair: Michel Waldschmidt

MA 005

08 Stephen Wainger, University of Wisconsin, Madison, USA Discrete Analogues of Singular and Maximal Radon Transforms (I: p.109) Chair: Detlef H. Müller

MA 001

09 Grzegorz Świa̧tek, Pennsylvania State University, University Park, USA Induced Hyperbolicity for One-Dimensional Maps (I: p.119) Chair: Klaus R. Schneider

H 1012

15 Hisashi Okamoto, RIMS, Kyoto University, Kyoto, Japan Bifurcation of Kolmogorov flows with an emphasis on singular limit (I: p.180) Chair: Rolf Jeltsch

H 1028

16 Gérard M. Iooss, Institut Universitaire de France, Valbonne, France Travelling water-waves, as a paradigm for bifurcations in reversible infinite dimensional “dynamical” systems (I: p.184) Chair: Klaus Kirchgässner

H 2032

19 Joseph Dauben, The City University of New York, New York, USA Marx, Mao and Mathematics: the Story of Nonstandard Analysis in China and the Cultural Revolution (I: p.208) Chair: Ivor Grattan-Guiness

MA 004

\underline{17:15 – 18:00}

03 William Duke, Rutgers University, New Brunswick, USA Bounds for arithmetic multiplicities (I: p.50) Chair: Per Salberger

MA 005

08 Vitali Milman, Tel Aviv University, Tel Aviv, Israel Randomness and Pattern in Convex Geometric Analysis (I: p.101) Chair: Hermann König

MA 001

09 Sergei B. Kuksin, Steklov Institute, Moscow, Russia and Heriot-Watt University, Edinburgh, UK Towards a Qualitative Theory of Hamiltonian PDEs (I: p.116) Chair: Jürgen Moser

H 1012

15 Gregory Beylkin, University of Colorado at Boulder, Boulder, USA On multiresolution methods in numerical analysis (I: p.178) Chair: Siegfried Prößdorf

H 1028

16 Ulf Grenander, Brown University, Providence, USA Strategies for Seeing (I: p.183) Chair: Friedrich Götze

H 2032

18 Mogens Niss, Roskilde University, Roskilde, Denmark Aspects of the nature and the state of research in mathematical education (I: p.198) Chair: Urs Kirchgraber MA 004

Short Communications and Poster Sessions

Short Communications: Monday, 24 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 03

Lattices and Codes, Chair: Fei Xu

MA 311 San Ling Group algebras and codes (p.41)

Katherine S. Roegner

Computing even, unimodular lattices with complete root systems (p.46)

Fei Xu

Arithmetic Springer Theorem on Quadratic Forms under Field Extensions of Odd Degree (p.48)

Sect.: 04

Moduli Spaces and Mathematical Physics, Chair: Ragni Piene

MA 042 Sheldon H. Katz Geometric Engineering of Quantum Field Theories (p.55)

Alexandre Kabanov

Intersection theory on the moduli spaces of stable curves (p.54)

Tyler J. Jarvis

Geometry of the Moduli of Higher Spin Curves (p.54)

Sect.: 07

Representations I, Chair: Carina Boyallian

H 112 Jose I. Liberati Quasifinite Highest Weight Modules over the Lie Algebra of Matrix Differential Operators on the Circle (p.107)

María L. Barberis

Quaternionic representations (p.101)

Paolo Casati

α–stratified modules of real reductive groups (p.104)

Sect.: 08

Several Complex Variables, Chair: Stefan B.A. Jakobsson

MA 041 Yoshihiro Abe Universal functions in several variables

(p.112)

Elias Wegert

Analytic discs attached to totally real surfaces (p.163)

Gabriela Kohr

Univalent mappings of several complex variables (p.135)

Sect.: 08

Real Analysis, Chair: Gilbert Helmberg

MA 309 Jan Hejcman A description of quasiconvex functions

(p.129)

Attila Gilányi

On polynomial functions, monomial functions, and the Dinghas interval derivative

(p.125)

Joris van der Hoeven

Proof of a conjecture of Hardy about Lfunctions (p.161)

Sect.: 08

Approximation Theory, Chair: Miguel A. Jimenez

MA 043 Alexander Yakovlevich Radyno The degree of rational approximation of a holomorphic function (p.149)

Yuri Kryakin

Polynomial Interpolation and Whitney’s Constants (p.138)

Miguel A. Jimenez

The Degree of Best Approximation in the Lipschitz Norm by Trigonometric Polynomials (p.133)

Sect.: 09

Singularly Perturbed Systems, Chair: Vladimir Sobolev

H 111 Minoru Nakano On the WKB method and Airy functions

(p.184)

Carmen Rocsoreanu

Canards in the Fitz-Hugh-Nagumo model (p.188)

Vladimir Sobolev

Canards and black swans (p.192)

Sect.: 09

Spectral Problems and Smooth Linearization, Chair: Vjacheslav A. Yurko

H 110 Shunian Zhang Stability of neutral delay difference systems (p.197)

Samovol Vladimir

The problem of local smooth linearization of systems of ordinary differential equations (p.189)

Vjacheslav A. Yurko

Inverse spectral problems for ordinary differential equations (p.196)

Sect.: 11

Geometry, General Relativity, Chair: Akira Asada

H 107 Mohamed Abdel-Megied Conformal Singularities (p.227)

Paula B. Cohen

Siegel’s Lemma, Padé Approximations and Jacobians (p.230)

Akira Asada

Clifford bundles over mapping spaces (p.228)

Sect.: 11

Thermodynamics, Hydrodynamics, Non-linear Dynamics II, Chair: Burkhard J. Schmitt

H 1029 Alexander Zlotnik Global weak solutions to the equations of one-dimensional motion of a viscous heat-conducting gas with nonsmooth data

(p.247)

Yoshimitsu Iwasaki

Axiomatic theory of the multiple anelastic relaxation in crystalline solids (p.235)

George Jaiani

A Mathematical Model of Bars with Variable Rectangular Cross-section (p.236)

Sect.: 15

Inverse Problems I, Chair: Anatoly G. Yagola

H 2033 Gottfried Bruckner On the regularization of an inverse vibration problem with noisy data (p.297)

Cesar Palencia

A numerical method for the integration of backward parabolic problems (p.308)

Sergii Pereverziev

An adaptive discretization of third kind Volterra equations with weakly singular kernels (p.309)

Sect.: 16

Control of Processes and Systems, Chair: Mihai M. Vladimirescu

H 106 Juergen Borchardt Large-scale Dynamic Process Simulation of Chemical Plants (p.320)

Elena I. Lyashenko

Optimal control of biotechnological process (p.328)

Mihai M. Vladimirescu

Rotative Actuators Robustness (p.333)

Sect.: 18

Pedagogical and Psychological Aspects, Chair: Hans-Joachim Arnold

H 2013 Sou-yung Chiu Pedagogical Mathematics Knowledge, PMK and P(PMK)K (p.358)

Torsten Fritzlar

Über Möglichkeiten der Sensibilisierung für die Komplexität von Mathematikunterricht (p.359)

Hans-Joachim Arnold

Geometrische Relationen-Algebra und Piaget’sche Schematheorie (p.357)

Short Communications: Monday, 24 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 03

Diophantine Problems, Chair: Pieter Moree

MA 311 Arturas Dubickas Polynomials vanishing at 1 with high multiplicity and Tarry’s problem (p.35)

Florian G. Luca

Equations Involving the Euler ϕ Function (p.42)

Pieter Moree

Distribution of primes having a prescribed primitive root over residue classes (p.43)

Sect.: 04

Vector Bundles II, Chair: Herbert Kurke

MA 042 Hagen Meltzer Exceptional vector bundles (p.57)

Vasile P. Brinzǎnescu

Moduli spaces of stable 2-vector bundles over ruled surfaces (p.51)

Elham Izadi

An embedding of a moduli space of vector bundles (p.54)

Sect.: 07

Representations II, Chair: Jose I. Liberati

H 112 Leandro R. Cagliero (K, M)-spherical submodules of the universal enveloping algebra of a real rank one semisimple Lie group G_{∘} (p.103)

Carina Boyallian

On Bernstein polynomials for compact Lie groups (p.103)

Tuong Ton-That

I Invariant Theory for some Infinite-Dimensional Lie Groups (p.110)

Sect.: 08

Harmonic Functions, Chair: Elias Wegert

MA 041 Anna Łazińska On a class of mappings harmonic in the half-plane, with a normalization at infinity

(p.138)

Promarz M. Tamrazov

The solution of Eremenko’s extremal problem on harmonic functions (p.159)

Eleutherius E. Symeonidis

The Poisson integral as an appropriate mean value (p.159)

Sect.: 08

A New View of Classical Phenomena, Chair: Gilbert Helmberg

MA 309 Andreas M. Hinz Catenaria vera — the true catenary

(p.130)

J. Marshall Ash

Nonuniqueness for smooth uniformly bounded orthonormal systems

(p.114)

Gilbert Helmberg

A corner-point Gibbs phenomenon for Fourier series in two dimensions (p.129)

Sect.: 08

Vector Lattices, Chair: Jacobus J. Grobler

MA 043 Rodica Mihaela Danet The extension of some linear operators

(p.121)

Nicolae Danet

Subspaces in directed vector spaces (p.121)

Jacobus J. Grobler

The polar decomposition of disjointness preserving operators (p.127)

Sect.: 09

Singularly Perturbed and Control Problems, Chair: Nikolay N. Nefedov

H 111 Klaus R. Schneider Exchange of stabilities in singularly perturbed systems (p.190)

Alexander G. Chentsov

Nonanticipating set-valued mappings and constructions of the programmed iterations method (p.174)

Nikolay N. Nefedov

Asymptotic method of differential inequalities for nonlinear singularly perturbed PDE’s with internal layers (p.185)

Sect.: 09

Stability of Dynamical Systems, Chair: Shaobo Gan

H 110 Abdelkader Boucherif A free boundary problem for a reactiondiffusion equation (p.172)

Juan H. Arredondo

A geometric approach to the asymptotic of systems of evolution equations (p.171)

Shaobo Gan

Star Systems X^{∗} and a Proof of the C^{1} Ω Stability (p.175)

Sect.: 11

Symmetries and Mathematical Methods in Physics, Chair: Vladimir V. Kisil

H 107 C. Martha Banulescu Lie Groups in Physics (p.229)

Malgorzata Klimek

Conserved Currents for Linear Equations of Motion in Discrete and Non-commutative Models (p.237)

Vladimir V. Kisil

Feynman Path Integral, Umbral Claculus, and Non-Commutative Combinatorics

(p.237)

Sect.: 11

Thermodynamics, Hydrodynamics, Non-linear Dynamics III, Chair: Alexander Zlotnik

H 1029 Burkhard J. Schmitt On upper bounds for the energy dissipation rate in turbulent shear flow (p.244)

Lev L.D. Pustylnikov

Poincare models, rigorous justification of the second law of thermodynamics on the basis of mechanics, and the Fermi acceleration mechanism (p.243)

Andrey B. Morgulis

����������� ��������������� ��������˘� ��������˘� � �������, �������� ��˘� ������ ����� �������˘� ����� (p.241)

Sect.: 15

Inverse Problems II, Chair: Sergii Pereverziev

H 2033 Sergiy Georgievich Solodkyy Optimization of projection methods for solving linear ill-posed problems (p.314)

Anatoly G. Yagola

Ill-posed problems and a priori information (p.318)

Weichung Wang

Using Preconditioned Conjugate Gradients in Final Iterations of Interior Point Methods (p.317)

Sect.: 18

Miscellaneous I, Chair: Masao Mori

H 2013 Salem A. Sahab Mathematics basic skills for grades 1 to 3

(p.363)

Josef Nietzsch

The Study of Mathematics by High School Students at the University (p.363)

Masao Mori

Beginners Confusion and Terminology in Mathematics (p.362)

Short Communications: Monday, 24 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 01

Constructive Mathematics and Models of Arithmetic, Chair: Hiloshi Ohno

MA 042 Douglas S. Bridges Ideals in constructive ring and Banach algebra theory (p.6)

Luminita Simona Dediu

A constructive extension theorem for ultraweakly continuous linear functionals (p.6)

Hiloshi Ohno

A certain proposition unprovable in number-theoretic formal system (p.9)

Sect.: 02

Algebras II, Chair: M. Susan Montgomery

MA 311 Mihai I. Cipu Multiplicities and pure resolutions for Gorenstein codimension four algebras

(p.15)

Michael Nüsken

Unique Tensor Factorization of Algebras (p.26)

Andrej A. Zolotykh

Standard bases for ideals of free algebras over commutative rings (p.32)

Sect.: 04

Singularities II, Chair: Herbert Kurke

MA 309 Michał Kwieciński An algebraic map with nonconstructible topology

Sect.: 05

Integrable Geodesic Flows and Perturbation Problems, Chair: Vladimir S. Matveev

MA 041 Masayuki Igarashi On Hermitian-Liouville structures on compact Hermitian surfaces (p.71)

Li Ma

Stable critical point Theory in perturbation problems (p.75)

Vladimir S. Matveev

Geodesic equivalence and integrability (p.76)

Sect.: 05

Local Differential Geometry, Chair: Jose Miguel Veloso

MA 043 Nikolaus K. Stephanidis Über einige Invarianten in der Theorie der Strahlensysteme (p.81)

Katerina Polyakova

Parallel displacements along manifold of osculating planes of a surface (p.78)

Jose Miguel Veloso

Almost pseudo-Hermitian manifolds (p.83)

Sect.: 06

Manifolds and Dynamics, Chair: Tadayoshi Mizutani

H 112 Levan A. Alania De Rham cohomology with local coefficients (p.85)

Alexander Fel’shtyn

Reidemeister torsion and dynamical zeta functions in Nielsen theory (p.88)

Efthimios Kappos

Manifolds: Differential Topology: Controllability of Vector Fields (p.90)

Sect.: 10

Elliptic Equations IV (Non-linear Problems), Chair: Hwai-Chiuan Wang

H 111 Carsten Ebmeyer The p-Laplacian under mixed boundary value conditions in multidimensional polyhedral domains (p.203)

Stanislav I. Pohozaev

Positive solutions of nonlinear Neumann problems. Application of the global fibering method (p.217)

Hwai-Chiuan Wang

Semilinear Elliptic Problems in Interior and Exterior Flask Domains (p.224)

Sect.: 10

Applied Problems I, Chair: Olga S. Rozanova

H 110 Zhiming Chen Adaptive Finite Element Methods for Dynamical GL Vortices in Superconductivity

(p.200)

Olga S. Rozanova

2-D model of atmosphere dynamics: energy balance and some properties of solutions (p.219)

Gregory A. Chechkin

On Behavior of Perforated Bodies with Rapidly Oscillating Surfaces

Sect.: 12

Liapunov Exponents, Chair: Michael K.R. Scheutzow

H 107 Volker Wihstutz Stability Maps for Systems with Noise Induced Stability (p.272)

Peter Imkeller

An explicit description of the Lyapunov exponents of the noisy damped harmonic oscillator (p.254)

Michel L.E. Rieutord

A boundary value problem of mixed type in a spherical shell (p.268)

Sect.: 12

Empirical Processes and Nonparametric Inference, Chair: Michael Nussbaum

H 1029 Samir Ben Hariz Uniform CLT for empirical process indexed by classes of functions (p.248)

Michael Jaehnisch

Asymptotic equivalence for a density model with independent but non identically distributed random variables (p.255)

Vladimir I. Koltchinskii

Empirical geometry of multivariate data: a deconvolution approach (p.257)

Sect.: 13

Classical Combinatorial Structures, Chair: Stefan Felsner

H 106 Mario-Osvin Pavcevic On two series of symmetric designs

(p.283)

Christian Krattenthaler

New results in the enumeration of rhombus tilings (p.280)

Martin Kochol

Superposition — a method for constructing graphs without nowhere-zero k-flows

(p.280)

Sect.: 17

0/1 Problems, Chair: Jannet A. Tambieva

H 2033 Friedrich Eisenbrand On the Chvátal rank of polytopes in the 0/1 cube (p.339)

Anzor D. Kasaev

Assigment Problem in Consideration of Discounting (p.344)

Jannet A. Tambieva

Research of 4-Cliques Problem (p.351)

Sect.: 18

Miscellaneous II, Chair: Mary E. Morley

H 2013 Ludwig Paditz Comparision of the statistic calculators TI-83, EL-9600, and CFX-9850G PLUS

(p.363)

Mary E. Morley

Three Response Types for Broadening the Conception of Mathematical Problem Solving in Computer-Graded Tests (p.362)

Zhongling Xu

The Popularization of Mathematical Methods in China (p.364)

Short Communications: Monday, 24 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 01

Recursion Theory and Model Theory, Chair: Helmut Pfeiffer

MA 042 David Hartz Hierarchy Based on Inductive Definitions

(p.7)

Cyrus F. Nourani

Functorial Computability And Generic Definable Models (p.8)

Sect.: 02

Hopf Algebras, Chair: Alexander A. Mikhalev

MA 311 Sorin Dascalescu Pointed Hopf algebras by Ore extensions and classification results (p.16)

M. Susan Montgomery

Prime Ideals in Hopf Galois Extensions (p.24)

Peter Schauenburg

Hopf algebras with a weak projection (p.27)

Sect.: 05

Homogeneous Spaces, Chair: Friedbert Prüfer

MA 041 Takashi Koda Weakly symmetric property of Riemannian 3-symmetric spaces (p.72)

Jan Slovák

Parabolic Geometries (p.80)

Friedbert Prüfer

Local homogeneity and differential operators (p.79)

Sect.: 05

Special Submanifolds/Exterior Differentiation, Chair: Peter Rögen

MA 043 Iaroslav A. Kopylov Normal solvability of the operator of exterior derivation on some surfaces in ℝ^{N}

(p.73)

Yoshio Matsuyama

On totally real submanifolds in CP^{n}(c) (p.76)

Peter Rögen

Embedding and knotting of flat surfaces in 3-space (p.79)

Sect.: 06

4-Manifolds, Chair: Alexander D. Mednykh

H 112 J. Scott Carter The Diagrammatics of Hopf Categories and Invariants of 4-Dimensional Manifolds

(p.86)

Nikolai Gusevskii

Complex hyperbolic surfaces and discrete groups in complex hyperbolic geometry

(p.88)

Alexander Prishlyak

Morse function conjugation on 4-manifolds (p.95)

Sect.: 10

Hyperbolic Equations III (Conservation Laws & Asymptotic Behavior), Chair: Armen Shirikyan

H 111 Christian Klingenberg Conservation laws with source terms

(p.209)

Andrzej Rogowski

Existence of space periodic solutions for semi-linear hyperbolic systems (p.218)

Armen Shirikyan

Asymptotic behaviour of solutions to second-order hyperbolic equations (p.221)

Sect.: 10

Reaction-Diffusion Equations, Chair: Klaus Zacharias

H 110 Gideon A. Ngwa The Effects of Boundary Conditions on Reaction Diffusion Equations (p.216)

Arnd Scheel

Essential instability of pulses amd bifurcations to modulated travelling waves

(p.220)

Klaus Zacharias

On a reaction–diffusion system modelling chemotaxis (p.225)

Sect.: 12

Large Deviations, Chair: Dmitry Ioffe

H 107 Esa Nummelin Large deviations of random vector fields with applications to economics (p.264)

Rolando Rebolledo

Stopping Times and Large-Deviation Principles (p.267)

Tzuu-Shuh Chiang

Large deviation of small diffusions in discontinuous media (p.250)

Sect.: 12

Statistics for Stochastic Differential Equations, Chair: Mikhailo Yadrenko

H 1029 Marianne Huebner Nonparametric estimation for stochastic partial differential equations Ruth Pfeiffer Statistical Problems for Stochastic Processes with Hysteresis (p.266)

Alexander Frolovich Taraskin

On problems of statistics for SDE (p.270)

Sect.: 13

Matrices, Posets, and Lattices, Chair: Jens Gustedt

H 106 Konrad Engel An upper bound for the size of a largest antichain in the poset of partitions of an integer (p.276)

Antoni Smoluk

Fix-point property of finite partially ordered sets (p.284)

Christian Bey

An intersection theorem for the function lattice (p.274)

Sect.: 17

Algorithms and Complexity Issues, Chair: Naum Zuselevich Shor

H 2033 Vitaly A. Perepelitsa Fractal Graphs and their Properties

(p.347)

Elena V. Popova

Estimates of Complexity of the Vector Ordering Problem (p.348)

Naum Zuselevich Shor

Methods of improving dual bounds with the use of functionally superfluous constraints in polynomial and boolean problems (p.350)

Sect.: 19

Non-European Mathematics, Chair: Shirong Guo

H 2013 Shirong Guo Yang Hui’s Influence on the Mathematical Tradition of the Ming Dynasty in China

(p.367)

Osamu Takenouchi

Summation of a sequence of natural numbers (p.369)

Abdelkhalek Cheddadi

Un problème d’arithmétique et sa résolution par le mathématicien marocain al-Hassar (XIIè s.) (p.365)

Poster Session: Monday, 24 August, 14:00–15:45

Sect.: 01  Logic

MA 144 Fariborz Azarpanah Sum and intersection of summand ideals in C(X) (p.5)

Grzegorz Bancerek

Mizar — a work environment for a mathematician: an experiment with continuous lattices (p.6)

Peter Michael Schuster

A Sequence Model for Constructive Nonstandard Analysis (p.10)

Sect.: 06  Topology

MA 144 Laura M. Anderson Matroid bundles and characteristic classes

(p.85)

Lorena Armas-Sanabria

3-Pure Braids and Surgery (p.85)

Vitali Chatyrko

On infinite-dimensional Cantor manifolds (p.86)

Mario Eudave-Munoz Incompressible surfaces in the exterior of tunnel number one knots (p.88)

Norikazu Hashiguchi

The surgery formula for the discrete Godbillon-Vey invariant (p.89)

Thang T.Q. Le

On Quantum and Finite Type Invariants of Homology 3-Spheres (p.91)

Jouni Luukkainen The fundamentals of Lipschitz and quasiconformal manifolds (p.92)

Leonard Mdzinarishvili

Partially continuous Alexander-Spanier cohomology theory (p.93)

Alexey Rasskazov

The structure of the canonical fundamental set for the 2-bridge link (p.95)

Sascha A. Rogmann Growth functions of tessellations (p.96)

Richard J. Steiner

Pasting in multiple categories (p.98)

Paul R. Turner

The homology of spaces representing exact cohomology theories (p.99)

Poster Session: Monday, 24 August, 16:15–18:00

Sect.: 07

Lie Groups and Lie Algebras

MA 144 Helge Glöckner Scale functions on p-adic Lie groups

(p.105)

Nikolai Gromov

Nonsemisimple Cayley-Klein groups and its quantization (p.105)

Duncan J. Melville

A survey of quantum deformation results (p.108)

Victor M. Petrogradsky On subalgebras of free Lie algebras

(p.108)

Jorge Soto-Andrade

Gel’fand Models, Weil Representations, and Twisted Spherical Functions (p.110)

Sect.: 11

Mathematical Physics

MA 144 Jack Asavanant Free-surface flows past a rectilinear object with stagnation points (p.228)

Valery Dryuma

Integration of Einstein equation for higherdimensional space (p.232)

Nadja Kutz

Discrete Nonlinear Toda Evolution (p.239)

David Natroshvili Direct and inverse problems of fluidstructure interaction (p.241)

Jean Ndimubandi

The q-Dirac oscillator: the context of radial potentials (p.242)

Joachim Rehberg

A hybrid semiconductor device model including nanostructures (p.243)

Ryuichi Sawae An action of Virasoro algebra on the formal loop group solutions of 2d reduced Einstein equations (p.244)

Joseph C. Várilly

Quantum fields over noncommutative tori (p.246)

Special Activities

\underline{14:00 – 18:00 H 104}

Session on Electronic Publishing Chair: Peter Michor, Vienna, Austria

14:00 – 14:30 Jane Kister, Mathematical Reviews Bernd Wegner, Zentralblatt für Mathematik The year 2000 Mathematics Subject Classification, MSC2000

14:30 – 15:00 Judith Plümer, Universität Osnabrück, Germany Mathematical Metadata

15:00 – 15:30 Stephen Buswell, Stilo Technology, UK An Introduction to MATHML

15:30 – 15:45 Discussion
15:45 – 16:15 Coffee break

16:15 – 18:00 Panel Discussion The future of electronic communication, information and publishing Panelists: N. Calkin, J. Ewing, M. Grötschel, J. Heinze, C. Hutchins, P. Michor, R. Palais Chair: David Mumford, IMU president

Other Events

\underline{17:30 Urania}

Orm Finnendahl, Berlin, Germany Konstruktion und musikalischer Sinn: Auswirkungen rekursiver und stochastischer Verfahren in meinen Kompositionen (in German)

\underline{19:30 Urania}

Hans–Magnus Enzensberger, München, Germany Zugbrücke außer Betrieb, oder die Mathematik im Jenseits der Kultur — eine Außenansicht (in German)